Timothys age in 15 years will be three times what it was 13 years ago how old is Timothy now
Answer by Timnewman(323)

Hi dear,
let Timothy's age be x
In fifteen years,
He will be (15+x)yrs
13yrs ago,he was (x-13)yrs
If his age in 15yrs is 3times his age 13yrs ago,
Then,
15+x=3(x-13)
15+x=3x-39
3x-x=15+39
2x=54
x=27
Timothy's age now is 27yrs old
HOPE THIS HELPS?
Best wishes,
Timnewman::